# Analytical coarse grained potential parameterization by Reinforcement Learning for anisotropic cellulose

### v2

@@ 3.4.1 Draw-out and Tear-apart
Figure 8: Draw-out and Tear-apart. (a) Draw-out and (b) curves. (c) Tear-apart and (d) curves. The transverse stretch resembled an opening fracture, whereas the in-plane and out-of-plane shear were examined using draw-out and tear-apart simulations. The draw-out and tear-apart were performed separately under the NPT and NVT ensembles by steered molecular dynamics with axial towing of the middle cellulose chain or perpendicular towing of the middle chains. The force-displacement curve of the draw-out presented periodic fading force curves that were ruled by their structure, whereas the tear-apart force was incremental. The maximum force and frequency of the draw-out and tear-apart matched those of the AA references. Therefore, the CG model can quantitatively represent the in-plane and out-of-plane shear.
@@ 3.4.2 Adhesion and Bending
Although our CG model was designed for pure periodic systems, adhesion and bending loads were applied to verify its applicability to aperiodic systems (Figure 9). An adhesion test was performed to determine the adhesion energy of nearby aperiodic CNC in the axial direction. A vertical load was applied at the center point of the cellulose bundle for the bending test. The adhesion energy of the CG model was evidently larger than that of AA, and the slightly lower force of the CG model in bending may result from the BD force constant rescaling of the angles and dihedrals. The results confirmed that the CG model could also be used in aperiodic CNC systems with slightly lower accuracy.
@@
Figure 9: Adhesion and Bending in aperiodic systems. (a) Adhesion. Two CNC were placed tightly close along the chain direction and were clung by the interface adhesion energy. They were then torn away from one another. The adhesion energy of the CG model was larger. (b) Bending. A vertical load was applied at the center of the CNC bundle to induce bending. The force curves of adhesion and bending confirm the capabilities of the CG model in aperiodic systems. Unlike the axial modulus dominated by bonds, the bending is also affected by the angles and dihedrals. The slightly lower bending modulus was assumed to be caused by the rescaling of the angle and dihedral force constants.
@@ 3.4.3 Transverse arrangement
Unlike previous BD directionality CG models[36, 38, 37, 39], the CG model was designed to present directionality via branched beads and HBond-like interactions. This makes it possible to consider the transverse arrangements of CNC with cross and symmetric patterns. The results of the arrangement patterns confirmed that the fractures and performance differed significantly from those of the crystals, as shown in Figure 10. The symmetric pattern was inspired by the twin structure, and the results indicated that this pattern exhibited a better toughness than the crystal. The cross pattern was symmetric in both directions. In these two patterns, frictional sliding was impeded and transformed into local rotations, thereby achieving better toughness and ductility in a symmetric pattern. These results emphasize the importance of the transverse arrangement, which cannot be described by former coarse-grained models.
@@ 3.4.4 Brick-and-mortar
Figure 11: Brick-and-mortar. (a) Brick-and-mortar models and fractures. (b) Stress-strain curves when the elementary cellulose chain length was 40 residues. (c) Strength and (d) toughness curves of different elementary chain lengths. Highly aligned CNC can assemble into brick-and-mortar structures [74, 75] in the axial direction, and this staggered structure can improve strength and toughness. The brick-and-mortar stretch was also a CNC-level in-plane shear, and was heavily influenced by the length of the elementary cellulose chains. The CG model accurately describes the structural behavior of the brick-and-mortar. The brick-and-mortar fracture was composed of major local slips and decentralized minor slips, and the toughness is positively correlated with the overlapping length.
@@ 3.4.5 Computational efficiency
The computational efficiencies of the AA and CG models were compared using equilibrium simulations. A sequence of square periodic CNC (the word square indicates that the row and column numbers of the transverse section of the CNC were the same) equilibrium simulations of different sizes were used, in which the atom number of the AA models ranged from seven kilos to one million. The benchmark simulations were performed using central processing units (CPU) and graphical processing units (GPU). The HB interaction does not support hardware acceleration, and the AA simulations using only the same amount of CPU resources are shown in Figure 12. For the AA and CG simulations using only the CPU, six cores were allocated exclusively to each task. Six cores and an additional exclusive GPU were allocated for the accelerated AA simulations using a GPU. The results confirmed that the CG model was 20 times faster than its AA counterpart when using the same resources. Even when referring to the AA models with more acceleration, the efficiencies CG models were triple. When a GPU-accelerated HB interaction is implemented and tested in the future, its potential can be fully exploited.
Figure 12: Computation efficiency data (a) and (b) relative efficiency data with AA simulations as a baseline. Currently, the HB interactions used does not support hardware acceleration. Referring to the AA simulations with same amount of CPU resources, the CG model was 20 times faster in CNC equilibrium simulations of different transverse sizes. The efficiency of the CG model was still triple when additional exclusive GPU resources were utilized for AA. However, its computational efficiency has not yet been fully proven.
@@
Cellulose materials are usually composed of long and thin cellulose nanofibrils, and previous CG models of cellulose have emphasized their axial properties[29, 30]. However, CNC are anisotropic and much shorter, and their transverse properties should not be disparaged, which is important for future development of cellulose materials. Some exhibit transverse anisotropy using anisotropic particles[26, 35] or additional artificial bonds[36, 37, 38, 39], which are still incapable of representing the transverse anisotropy.
@@
The challenges and main characteristics of the CG model include the flat residues and HBond (HBond) interactions. The CG model could reveal the frictional sliding and transverse arrangement, which are two critical properties for CNCs transverse anisotropy, and the former CG models could not reflect accurately. Except for the HBond interactions, only the harmonic and Lennard-Jones potentials were included in the BD and NB interactions. This model also confirms the potential parameterization capability of RL; only a discriminant reward function was required. The reward function for RL includes the matching degree of the BD properties (axial modulus and polymer stiffness) and the NB properties (transverse strength and toughness). The resultant model can reproduce the xyz three-dimensional anisotropy and transverse frictional sliding. RL may be more widely applied to potential parameterizations in the future.
@@
However, several problems are encountered in the CG model. The first is he rescale of the angle and dihedral interaction force constants. The CG model, which has significantly fewer degrees of freedom, must compensate for the polymer stiffness properties by decreasing the force constants (Table SP2). The former CG model addressing persistence length did not require a precise match (40 nm)[34]. Some studies also estimated the persistence length of a bundle of cellulose chains instead of a single free chain[76]. These smaller force constants may cause problems for other properties involving the angles and dihedrals. For example, a lower stress was observed during the bending test.
@@
The nonbonded equilibrium distances are also problematic. The equilibrium distances of the CG model were larger than those of their AA counterparts (Figure SP9), and the enlargement of the CL2-CL3 NB distance coefficients was determined using RL (Figure 5). This means that the distance coefficients and equilibrium distances do not correspond to the well-relaxed AA equilibrium distances or radial distribution function. In this study, the equilibrium distance of a pair was defined as the average minimum distance. Nearly all minimum distances of pairs are from beads of nearby HBonding layers; only the minimum distances of CL2-CL3 are from beads in the same HBonding layer. Therefore, we infer that the equilibrium distances from nearby HBonding plane pairs are appropriate estimations of $\sigma$ values for the LJ potentials because the dominant NB interactions are HBonds in the HBonding layers. The overall enlargement of the NB distance coefficients is assumed to be required by the increased $\sigma_{p23}$ to retain the crystal structure. Moreover, the enlargement of $\sigma_{p23}$ may be an assimilation of $\sigma$ because the equilibrium distance between CL2 and CL3 is the smallest. The Phenomenon of frictional sliding may be presented more easily if all $\sigma$ values are close to each other.
@@
Other minor problems also exist. Our model may not perform well in aperiodic cellulose systems (as in the adhesion test). The pressure curve of the CG model under the NPT ensemble exhibited a larger oscillation amplitude. Otherwise, some molecular details were ignored in the CG model. For instance, the rotational degrees of freedom of the hydrogen atoms on the branch hydroxyls were neglected in harmonic BD potentials (Figure SP9). Moreover, one technical problem temporarily restrained the application potential of the CG model: the HBond-like pair style does not support hardware acceleration. Therefore, the computational efficiency of the CG model has not yet been fully realized.
@@
As an early bird that introduced HBonds, the CG model of CNC could promote further research. The development of cellulose CG models has mainly focused on the native cellulose. However, solvent environments, particularly humidity, are key factors for cellulose materials[34, 77, 78, 79] and are not well represented in cellulose CG models[30]. Chemical surface modifications and interactions with other molecules such as hemicellulose or lignin are also critical for enhancing the performance of cellulose materials. These problems are tightly connected to the hydroxyls of cellulose residues. Regarding the computational resources and expertise, these interactions were not considered in this study. With the help of HBond-like interactions, a later CG model could attempt to present more properties of the cellulose. We hope that this model will facilitate future development of CG models.
@@
As an application of RL for potential parameterization, our methods and implementations are not heavily dependent on cellulose, and may help in later CG model development. In the CG model, many priori knowledge including pre-defined mapping and interaction potential expressions are used to drop the difficulty and enhance physical explainability. However, the resultant effect of priori knowledge is not sure. Fully automatic optimization may achieve models with better performances. Transfer Learning and Imitation Learning methods may also be applied in the future to further accelerate RL in the potential parameterization applications.
@@ Conclusions
In this study, we introduce a novel coarse-grained cellulose model with explicit branch beads and HBond-like interactions that can reproduce the transverse frictional sliding and anisotropic mechanical properties of cellulose nanocrystals. The strength and toughness in the three characteristic directions of the orthotropic transverse section were quantitatively reproduced without anisotropic particles or additional bonds. This study is an early bird that applied degenerate Reinforcement Learning as a nonlinear optimizer to perform direct analytical potential parameterization, and the physical meanings were enhanced by integrating the Boltzmann inversion method. This model is eclectic of physical knowledge, traditional modeling, and Machine Learning methods. It assists in the understanding of the importance of hydrogen bonds and the transverse anisotropy. With an appropriate reward function based on the target properties and statistics-guided reduction, the Reinforcement Learning agent can determine the appropriate coefficients without consuming considerable computational resources. The target anisotropic properties considered included the bonded properties (axial elastic modulus and polymer stiffness) and nonbonded properties (transverse stretch strength and toughness), all of which were accurately matched with errors below %15. Several additional inspections were performed after the training, including in-plane shear, out-of-plane shear, adhesion energy, bending, transverse arrangement, and brick-and-mortar structures. The transferability and generalization of this coarse-grained model were proven through additional inspections. These results emphasize that transverse anisotropy is important for transverse arrangements, which has not been quantitatively presented in previous coarse-grained models. Two transverse arrangement patterns with improved toughness and ductility were introduced. We found a positive correlation between the overlapping length and toughness of the brick-and-mortar structures. Data on transverse arrangements and brick-and-mortar structures may aid advanced material design in the future. This model may also aid in addressing interactions between cellulose and other polar molecules on a mesoscopic scale. This study emphasizes that Machine Learning can also be applied to construct effective analytical potentials that are better in terms of explainability, transferability, and kinetics. Although our results demonstrate the efficiency of Reinforcement Learning, many refinements remain to be explored, especially in the field of molecular research. For example, the necessity of inspiration from human knowledge and traditional methods such as Boltzmann inversion is still unclear. Finally, our basic parameterization approach is agnostic to the application case, implying that our workflow may also be applicable to other molecular systems. Transfer Learning and Imitation Learning methods may also be integrated with Reinforcement Learning to solve similar problems in the future.
Code availability
The modified hbond/dreiding/lj/angleoffset was merged in the official LAMMPS at
https://github.com/lammps/lammps/tree/develop/src/EXTRA-MOLECULE.
The code for potential parameterization training is available at
https://github.com/EiPiFun/rl-cll-cg.
Training data for the trained coefficients, pre-processing codes, and post-processing codes are available upon requests.
@@ Acknowledgments
Xu Don thanks the Department of Engineering Mechanics of Zhejiang University for the funding and computational resources.
@@
Force and energy curve of HB and LJ potentials
All atom transverse stretch performance at different stretch speed
@@
Other structural properties
Secondary summits
Table SP1:
Table SP2:
Figure SP1:
Figure SP2:
Figure SP3:
Figure SP4:
Figure SP5:
Figure SP6:
Figure SP7:
Figure SP8:
Figure SP9:

### v3

@@ 3.7.1 Draw-out and Tear-apart
Figure 9: Draw-out and Tear-apart. (a) Draw-out and (b) curves. (c) Tear-apart and (d) curves. The transverse stretch resembled an opening fracture, whereas the in-plane and out-of-plane shear were examined using draw-out and tear-apart simulations, respectively. The draw-out and tear-apart were performed separately under the NPT and NVT ensembles by steered molecular dynamics with axial towing of the middle cellulose chain or perpendicular towing of the middle chains. The force-displacement curve of the draw-out presented periodic fading force curves that were ruled by their structure, whereas the tear-apart force was incremental. The maximum force and frequency of the draw-out and tear-apart matched those of the AA references. Therefore, the CG model can quantitatively represent the in-plane and out-of-plane shear.
@@ 3.7.2 Adhesion and Bending
Although our CG model was designed for pure periodic systems, adhesion and bending loads were applied to verify its applicability to aperiodic systems (Figure 10). An adhesion test was performed to determine the adhesion energy of nearby aperiodic CNC in the axial direction. A vertical load was applied at the center point of the cellulose bundle for the bending test. The adhesion energy of the CG model was evidently larger than that of AA, and the slightly lower force of the CG model in bending may result from the BD force constant rescaling of the angles and dihedrals. The results confirmed that the CG model could also be used in aperiodic CNC systems with a slightly lower accuracy.
@@
Figure 10: Adhesion and Bending in aperiodic systems. (a) Adhesion. Two CNCs were placed tightly close along the chain direction and were clung by the interface adhesion energy. They were then torn away from one another. The adhesion energy of the CG model was larger. (b) Bending. A vertical load was applied at the center of the CNC bundle to induce bending. The force curves of adhesion and bending confirm the capabilities of the CG model in aperiodic systems. Unlike the axial modulus dominated by bonds, the bending is also affected by the angles and dihedrals. The slightly lower bending modulus was assumed to be caused by the rescaling of the angle and dihedral force constants.
@@ 3.7.3 Transverse arrangement
Unlike previous BD directionality CG models[36, 38, 37, 39], the CG model was designed to present directionality via branched beads and HBond-like interactions. This makes it possible to consider the transverse arrangements of CNCs with cross and symmetric patterns. The results of the arrangement patterns confirmed that the fractures and performance differed significantly from those of the crystals, as shown in Figure 11. The symmetric pattern was inspired by the twin structure, and the results indicated that this pattern exhibited a better toughness than the crystal. The cross pattern was symmetric in both directions. In these two patterns, frictional sliding was impeded and transformed into local rotations, thereby achieving better toughness and ductility in a symmetric pattern. These results emphasize the importance of the transverse arrangement, which cannot be described by former CG models.
@@ 3.7.4 Brick-and-mortar
Figure 12: Brick-and-mortar. (a) Brick-and-mortar models and fractures. (b) Stress-strain curves when the elementary cellulose chain length was 40 residues. (c) Strength and (d) toughness curves of different elementary chain lengths. Highly aligned CNCs can assemble into brick-and-mortar structures [95, 96] in the axial direction, and this staggered structure can improve strength and toughness. The brick-and-mortar stretch was also a CNC-level in-plane shear, and was heavily influenced by the length of the elementary cellulose chains. The CG model accurately describes the structural behavior of the brick-and-mortar. The brick-and-mortar fracture was composed of major local slips and decentralized minor slips, and the toughness is positively correlated with the overlapping length.
@@ 3.7.5 Computational efficiency
The computational efficiencies of the AA and CG models were compared using equilibrium simulations. A sequence of square periodic CNCs (the word square indicates that the row and column numbers of the transverse section of the CNCs were the same) equilibrium simulations of different sizes were used, in which the atom number of the AA models ranged from seven kilos to one million. The benchmark simulations were performed using central processing units (CPU) and graphical processing units (GPU). The HB interaction does not support hardware acceleration, and the AA simulations using only the same amount of CPU resources are shown in Figure 13. For the AA and CG simulations using only the CPU, six cores were allocated exclusively to each task. Six cores and an additional exclusive GPU were allocated for the accelerated AA simulations using a GPU. The results confirmed that the CG model was 20 times faster than its AA counterpart when using the same resources. Even when referring to the AA models with additional GPU acceleration, the efficiencies of the CG models were triple. When a GPU-accelerated HB interaction is implemented and tested in the future, its potential can be fully exploited.
Figure 13: Computation efficiency. (a) efficiency data and (b) relative efficiency data with AA simulations as a baseline. Currently, the HB interactions used does not support hardware acceleration. Referring to the AA simulations with same amount of CPU resources, the CG model was 20 times faster in CNC equilibrium simulations of different transverse sizes. The efficiency of the CG model was still triple when additional exclusive GPU resources were utilized for AA. However, its computational efficiency has not yet been fully proven.
@@
Cellulose materials are usually composed of long and thin cellulose nanofibrils, and previous CG models of cellulose have emphasized their axial properties[29, 30]. However, CNCs are anisotropic and much shorter, and their transverse properties should not be disparaged, which is important for future development of cellulose materials. Some exhibit transverse anisotropy using anisotropic particles[26, 35] or additional artificial bonds[36, 37, 38, 39], which are still incapable of representing the transverse anisotropy.
@@
The challenges and main characteristics of the CG model include the flat residues and hydrogen bond (HBond) interactions. The CG model could reveal the laminar structures and frictional sliding, which are two critical properties for CNCs transverse anisotropy, and the former CG models could not reflect accurately. Except for the HBond interactions, only the harmonic and Lennard-Jones potentials were included in the BD and NB interactions. This model also confirms the potential parameterization capability of RL; only a discriminant reward function was required. The reward function for RL includes the matching degree of the BD properties (axial modulus and polymer stiffness) and the NB properties (transverse strength and toughness). RL may be more widely applied to potential parameterizations in the future.
In the context of CG modeling, this model owns its uniqueness. In principle, it followed bottom-up approach starting from molecular details but used strength and toughness as optimization target, rather than potential of mean force or other structural distributions. And it chose simple analytical function form, which is also similar to top-down to a certain extent. Its mapping and topology restores molecular shape and frictions, and the interaction functions implemented anisotropic HBonding. And more importantly, it was designed to explore equilibrium states and processes involving dynamics properties with quantitative accuracy. On the other hand, this model preserved necessary molecular structures, and may be suitable for AA reconstruction based upon CG models.
@@
However, several problems are encountered in the CG model. The first is the rescale of the angle and dihedral interaction force constants. The CG model, which has significantly fewer degrees of freedom, must compensate for the polymer stiffness properties by decreasing the force constants (Table Supplementary Information). The former CG model addressing persistence length did not require a precise match (40 nm)[34]. Some studies also estimated the persistence length of a bundle of cellulose chains instead of a single free chain[97]. These smaller force constants may cause problems for other properties involving the angles and dihedrals. For example, a lower stress was observed during the bending test.
@@
The nonbonded equilibrium distances are also problematic. The equilibrium distances of the CG model were larger than those of their AA counterparts (Figure Supplementary Information, Figure Supplementary Information), and the enlargement of the CL2-CL3 NB distance coefficients was determined using RL (Figure 5). This means that the distance coefficients and equilibrium distances do not correspond to the well-relaxed AA equilibrium distances or radial distribution function. In this study, the equilibrium distance of a pair was defined as the average minimum distance. Nearly all minimum distances of pairs are from beads of nearby HBonding layers; only the minimum distances of CL2-CL3 are from beads in the same HBonding layer. Therefore, we infer that the equilibrium distances from nearby HBonding plane pairs are appropriate estimations of $\sigma$ values for the LJ potentials because the dominant NB interactions are HBonds in the HBonding layers. The overall enlargement of the NB distance coefficients is assumed to be required by the increased $\sigma_{p23}$ to retain the crystal structure. Moreover, the enlargement of $\sigma_{p23}$ may be an assimilation of $\sigma$ because the equilibrium distance between CL2 and CL3 is the smallest. The Phenomenon of frictional sliding may be presented more easily if all $\sigma$ values are close to each other. This also influenced the accuracy of $theta_{0}$ in the HB potential. In the first reduction, the distance coefficients between pairs of same kinds were artificially aligned to be the same because laminar structure and coefficients are critical for CG model. To align with the artifacts constraints, $theta_{0}$ in the HB potential was determined by the mapped artificial model, which was constructed by coordinate offset and ideally parallel. Thus, the $theta_{0}$ is an approximation without strong theory background.
@@
Other minor problems also exist. Our model may not perform well in aperiodic cellulose systems (as in the adhesion test). The pressure curve of the CG model under the NPT ensemble exhibited a larger oscillation amplitude. Otherwise, some molecular details were ignored in the CG model. For instance, the rotational degrees of freedom of the hydrogen atoms on the branch hydroxyls were neglected in harmonic BD potentials (Figure Supplementary Information). Moreover, one technical problem temporarily restrained the application potential of the CG model: the HBond-like pair style does not support hardware acceleration. Therefore, the computational efficiency of the CG model has not yet been fully realized.
@@
As a pioneering work that introduced HBonds, the CG model of CNC could promote further research. The development of cellulose CG models has mainly focused on the native cellulose. However, solvent environments, particularly humidity, are key factors for cellulose materials[34, 98, 99, 100] and are not well represented in cellulose CG models[30]. Chemical surface modifications and interactions with other molecules such as hemicellulose or lignin are also critical for enhancing the performance of cellulose materials. These problems are strongly dependent on the hydroxyls of cellulose residues. Regarding the computational resources and expertise, these interactions were not considered in this study. With the help of HBond-like interactions, a later CG model could attempt to present more properties of the cellulose. We hope that this model will facilitate future development of CG models.
@@
As an application of RL for potential parameterization, our methods and implementations are not heavily dependent on cellulose, and may help in later CG model development. In the CG model, much priori knowledge, including pre-defined mapping and interaction potential expressions, is used to reduce the difficulty and enhance physical explainability. However, the resultant effect of priori knowledge is uncertain. Fully automatic optimization may achieve models with better performances. Transfer Learning and Imitation Learning methods may also be applied in the future to further accelerate RL in the potential parameterization applications.
@@ Conclusions
In this study, we introduce a novel CG cellulose model with explicit branch beads and HBond-like interactions that can reproduce the anisotropy of cellulose nanocrystals. The strength and toughness in the three characteristic directions of the orthotropic transverse section were quantitatively reproduced without anisotropic particles or additional bonds. This study is a pioneering work that applied degenerate RL as a nonlinear optimizer to perform direct analytical potential parameterization, and the physical meanings were enhanced by integrating the Boltzmann inversion method. This model may also contribute to the CG modeling in its own way. Rather than targeting at traditional mapped structural distributions, this CG model extended bottom-up approach by implementing anisotropic interactions and hydrogen bonding and incorporating mechanical properties. The obvious advantage is the simple analytical function forms supported by optimized mapping and topology that explore dynamics properties quantitatively. It is a hybrid-hybrid model: combing bottom-up and top-down, considering both equilibrium and non-equilibrium. Such a not too coarse model may also benefit multiscale simulations which requires backmapping and reconstruction. This model synthesizes physical knowledge, traditional modeling, and Machine Learning methods. It aids in the understanding of the importance of hydrogen bonds and the transverse anisotropy. With an appropriate reward function based on the target properties and statistics-guided reduction, the RL agent can determine the appropriate coefficients without consuming considerable computational resources. The target anisotropic properties considered included the bonded properties (axial elastic modulus and polymer stiffness) and nonbonded properties (transverse stretch strength and toughness), all of which were accurately matched with errors below 15%. Several additional inspections were performed after the training, including in-plane shear, out-of-plane shear, adhesion energy, bending, transverse arrangement, and brick-and-mortar structures. The transferability and generalization of this CG model were proven through additional inspections. These results emphasize that transverse anisotropy is important for transverse arrangements, which has not been quantitatively captured in previous CG models. Two transverse arrangement patterns with improved toughness and ductility were introduced. We found a positive correlation between the overlapping length and toughness of the brick-and-mortar structures. Data on transverse arrangements and brick-and-mortar structures may aid advanced material design in the future. This model may also aid in addressing interactions between cellulose and other polar molecules on a mesoscopic scale. This study emphasizes that Machine Learning can also be applied to construct effective analytical potentials that are better in terms of explainability, transferability, and kinetics. Although our results demonstrate the efficiency of RL, many refinements remain to be explored, particularly for molecular systems. For example, the necessity of inspiration from human knowledge and traditional methods such as Boltzmann inversion is still unclear. Finally, our basic parameterization approach is general to the applications, implying that our workflow may also be applicable to other molecular systems. Transfer Learning and Imitation Learning methods may also be integrated with RL to solve similar problems in the future.
@@
Data availability
The modified hbond/dreiding/lj/angleoffset was merged into the official LAMMPS EXTRA-MOLECULE at
https://github.com/lammps/lammps
. The codes for potential parameterization training are available at
https://github.com/EiPiFun/rl-cll-cg
. The training data, models, simulation codes, processed data and plot scripts are available at
https://github.com/EiPiFun/rl-cll-cg-data
. The raw data, pre-processing codes, and post-processing codes are available upon requests.
@@ Acknowledgments
Xu Dong thanks the Department of Engineering Mechanics of Zhejiang University for the funding and computational resources.
Author Contributions
Xu Dong: Conceptualization, Data curation, Formal analysis, Investigation, Methodology, Project administration, Resources, Software, Validation, Visualization, Writing - original draft, Writing - review & editing.
A professor from the Department of Engineering Mechanics of Zhejiang University provided some additional Methodology, Resources, and Writing - review & editing helps and chose to be anonymous.
@@
Force and energy curve of HB and LJ potentials
Sensitivity and ablation check for potential formula
All atom transverse stretch performance at different stretch speed
Temperature and pressure curve using 12 fs timestep
Component match degree of training loops
@@
Transverse strength and toughness distributions
@@
Additional structural properties
Baseline method structures for coarse grained models
Training using reward function without threshold
Training using other optimization methods
Secondary summits of the bonded geometry parameter distributions
Table S1, Table S2, Table S3, Table S4
Figure S1, Figure S2, Figure S3, Figure S4, Figure S5, Figure S6, Figure S7, Figure S8,
Figure S9, Figure S10, Figure S11, Figure S12, Figure S13, Figure S14, Figure S15,
Figure S16, Figure S17, Figure S18, Figure S19, Figure S20